Solve 3(x-1)-(5x-5)=12
Found 2 solutions by FrankM, Tatiana_Stebko:
Answer by FrankM(1040)

3(x-1)-(5x-5)=12
3x-3-5x+5=12
-2x+2=12
-2x=10
x=10/(-2)
x=-5

Answer by Tatiana_Stebko(1539)